the sum of a number and twice the second number is 15. when the second number is subtracted from the first number, the difference is 3. find the two numbers
Answer by stanbon(75887)

the sum of a number and twice the second number is 15. when the second number is subtracted from the first number, the difference is 3. find the two numbers
------
Equations:
x + 2y = 15
x - y = 3
------
Subtract and solve for "y":
3y = 12
y = 4
-------
Solve for "x":
x-y = 3
x - 4 = 3
x = 7
==============
Cheers,
Stan H.
==============